### v1

# A deep learning model for chemical shieldings in molecular organic solids including anisotropy

Abstract
Nuclear Magnetic Resonance (NMR) chemical shifts are powerful probes of local atomic and electronic structure that can be used to resolve the structures of powdered or amorphous molecular solids. Chemical shift driven structure elucidation depends critically on accurate and fast predictions of chemical shieldings, and machine learning (ML) models for shielding predictions are increasingly used as scalable and efficient surrogates for demanding ab initio calculations. However, the prediction accuracies of current ML models still lag behind those of the DFT reference methods they approximate, especially for nuclei such as 13C and 15N. Here, we introduce ShiftML3, a deep-learning model that improves the accuracy of predictions of isotropic chemical shieldings in molecular solids, and does so while also predicting the full shielding tensor. On experimental benchmark sets, we find root-mean-squared errors with respect to experiment for ShiftML3 that approach those of DFT reference calculations, with RMSEs of 0.53 ppm for 1H, 2.4 ppm for 13C, and 7.2 ppm for 15N, compared to DFT values of 0.49 ppm, 2.3 ppm, and 5.8 ppm, respectively.
Chemical function is directly related to structure. As a result, the determination of atomic-level three-dimensional structures is at the heart of molecular and materials science. In this context, solid-state NMR (often in conjunction with diffraction methods and other spectroscopic methods) has emerged as the method of choice for atomic-level characterization of complex materials in powder form. This has been driven by the development of methods to probe the local atomic environment based on chemical shifts, [1, 2, 3, 4, 5, 6] often referred to as NMR crystallography. In turn, this has been made possible by the development of accurate density functional theory (DFT) based methods to calculate chemical shifts. [7, 8] Indeed, it has been recently shown that NMR can elucidate structures of molecular solids from powders with resolution better than 0.2 Å. [9, 10]
In a nutshell, NMR crystallography consists in experimentally measuring and assigning chemical shifts for each of the atomic sites in a structure, and then comparing the measured shifts to those predicted by computation for a pool of candidate structures, that are themselves enumerated by a computational search. Chemical shifts are traditionally calculated using DFT methods. [7, 11, 12, 13, 14] This workflow has been successful for materials ranging from hybrid-perovskites [15, 16] and inorganic materials [17, 18, 19, 20, 21, 22] to enzyme activate sites, [23, 24, 10] organic molecular solids, [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and notably for amorphous drug formulations. [40, 41, 42, 43] One of the key bottlenecks to address in this workflow is both the computational cost and the accuracy of chemical shift predictions.
The introduction of accurate machine learning models to predict chemical shifts in solids has essentially removed the computational cost barrier, allowing for tens of millions of shifts to be predicted on the order of days. [44, 45, 46, 47, 48, 49, 50, 51, 52, 53] For organic molecular solids, Paruzzo et al. [45] developed a model (dubbed ShiftML) that could predict isotropic chemical shifts for any molecular solid containing C, H, N, O atoms, that was trained on GIPAW DFT data for 2,000 structures from the Cambridge structural database (CSD). [54] Engel and coworkers later broadened its applicability to sulfur-containing compounds. [55] The model was subsequently improved and extended to ShiftML2 [49] which was trained on an extended set of over 14,000 structures containing H, C, N, O, S, F, P, Cl, Na, Ca, Mg and K, and composed of both relaxed and thermally perturbed structures. Both ShiftML models used kernel ridge regression (KRR), [56] with uncertainties in the predictions estimated using a committee of 32 models (for ShiftML2). [57] The local atomic environments were described using smooth overlap of atomic positions (SOAP) descriptors. [58]
The predictions of ShiftML2 yield an accuracy for isotropic 1H shifts of 0.5 ppm with respect to the DFT computed shifts, and an accuracy of 0.47 ppm with respect to a benchmark set of experimental shifts. [49] Increasing the accuracy of shift prediction for ML models would directly translate to better structure determination. Indeed, Unzueta et al. [48] have shown that using a $\Delta$-ML approach to improve the prediction accuracy of DFT computed shifts themselves, by correcting the effects of small basis set sizes in DFT shielding calculations, can lead to errors as small as 0.11 ppm for 1H shifts with respect to DFT. Although this is not directly relevant to pure ML predictions, since they use DFT computations as a starting point, it does nicely illustrate the potential to develop more accurate models.
Even though DFT calculations naturally yield the full chemical shielding tensor, the prediction models developed so far for molecular solids were only trained on isotropic values of the chemical shift (where the isotropic shift is the average of the three principal components). Recently, models for silicon oxides [51, 52] and non-magnetic oxides [53] have been introduced to predict the full chemical shift tensor with orientation, extending the scope of ML-predicted chemical shieldings to applications that rely on the chemical shift anisotropy (CSA). Tensorial models have also been used to predict other quantities relevant to nuclear spectroscopy, such as electric field gradients [59] and zero-field splitting tensors [60].
Here, we present a new machine learning model (dubbed ShiftML3) that predicts the full chemical shielding tensors, trained using shieldings calculated at the GIPAW/PBE level of theory. The model is based on an ensemble of Point Edge Transformer (PET) models. [61] The model not only predicts anisotropic shieldings, but also yields significantly improved accuracy with respect to DFT, with up to a two-fold improvement over ShiftML2 for the isotropic shieldings of heavy atoms.
Isotropic shifts are often sufficient to identify the experimentally-observed crystal structure out of a pool of potential candidates. For this reason, previous ShiftML models have been constructed by directly learning the isotropic chemical shieldings $\sigma_{\text{iso}}$ from DFT calculations, using symmetry-adapted models designed to make predictions that are exactly invariant to rigid rotations of a structure. However, the chemical shielding is a tensorial property, whose anisotropy (typically referred to as the chemical shift anisotropy or CSA) provides additional information that can be useful to explore crystal structures. [62] Given that GIPAW-DFT calculations yield the full chemical shielding tensor, in ShiftML3 we include tensorial predictions, focusing on the symmetric component which is of experimental relevance. For molecular solids, isotropic chemical shifts are readily measured using fast magic-angle spinning (MAS) methods, [4] and CSA can be measured using slow MAS [63] or using multi-dimensional isotropic-anisotropic correlation experiments to avoid spectral crowding. [64, 65, 66, 67] While these parameters are readily measurable, there is currently no sufficiently large database of experimental shifts for molecular solids with corresponding high-resolution atomic-level structures. This is in contrast to small molecules and proteins in solution, which have large and well-curated experimental databases. [68, 69] As a result, ML models in solids have been developed by using curated sets of structures together with shifts computed using DFT.
Here we use the training set from the most recent previous ShiftML model, and for which the computed shifts (computed at the PBE level of theory) contain the full tensor information. [49] This dataset is composed of molecular solids where the structures include (up to) 12 elements [H, C, N, O, F, Na, Mg, P, S, Cl, K, Ca] commonly found in organic compounds. It contains a total of 14,146 structures, including 7,548 sampled from the Cambridge structural database (CSD) via farthest point sampling (FPS) [70] to cover a broad range of structural motifs, and 6,598 additional thermally distorted structures. The ShiftML2 test set was divided into a new validation (505 structures, with 41,599 atomic environments) and a new test set (1,254 structures, with 106,230 atomic environments). In this split we choose a stratified selection strategy that ensures that both the distorted and the relaxed structure of any one CSD identifier are found either in the validation or the test set, avoiding data leakage between the sets. Furthermore, we ensure that both the new validation and new test sets contain at least a few environments of all 12 chemical elements. To compare the accuracies between ShiftML2 and ShiftML3 in a consistent fashion, the accuracies are reported on this new test set (referred to as the CSD-test set). More details of the training, validation and test sets for each element in CSD-test are given in section S1 of the supporting information. Due to computational constraints, ShiftML2 was trained on a representative subselection of atomic environments from the training set. The favorable computational scaling of neural network training via stochastic gradient descent allowed us to train ShiftML3 using all 1.4 million atomic environments contained in the 14,146 structures in the training set. The updated dataset containing the chemical shielding tensor data is available via a public data record associated with this publication. [71] We also use established datasets of experimental isotropical chemical shielding values, with 13 structures for 1H, 21 for 13C, and 15 for 15N, [49, 14] and a benchmark set of experimental 13C CSA principal values for 19 compounds. [72]
For the construction of the ShiftML3 model, we learn the correlation between chemical shielding tensor and structure using an ensemble of Point Edge Transformer (PET) models. [61] PET is an unconstrained graph neural network (GNN) architecture that delivers accurate and fast predictions for several material properties benchmark sets, importantly also on rotationally equivariant targets such as molecular dipole moments. [61] For ShiftML3, we use a modified version of the PET architecture referred to as nanoPET (available via the metatrain library). [73] Even though nanoPET does not enforce rotational invariance or equivariance of the target predictions, it learns the rotational behaviour of the targets via rotational augmentation during the training process – a mechanism that has been shown to yield models that are symmetry-compliant to a high accuracy. [74]
In practice, nanoPET predicts the chemical shielding tensors in their irreducible spherical tensor (IST) representation in a manner similar to symmetry-adapted, rotationally-constrained models. [75] The chemical shielding tensors are decomposed into their irreducible real spherical representations: A scalar, $\lambda=0$ part, a $\lambda=1$ pseudovector, and a $\lambda=2$ proper 5-vector having the same rotational behaviour as their corresponding $Y^{\lambda}$ spherical harmonics. [76] We refer the reader to the following references, [77, 78, 79] for a more in depth discussion of the properties of the chemical shielding tensor and a discussion of its rotational properties. nanoPET uses a single message-passing sequence of unconstrained graph transformer modules, with one final linear readout layer per IST component, $T^{\lambda}$. The final tensor predictions are then obtained by a change of representation back from the predicted IST components to the tensor in the Cartesian basis.
During the training procedure, random rotations in their irreducible Wigner-D matrix representations are automatically applied to the target IST components at each training step, as well as random inversions. We train an ensemble consisting of 8 nanoPET models. The models are trained for 1,500 epochs, with a constant learning rate of $3\cdot 10^{-4}$. We train the model using mini batch gradient descent with the Adam optimizer [80] and a batch size of 4. The construction of an ensemble model has several benefits; in addition to increased prediction accuracies, an ensemble also provides uncertainty estimates and further reduces the rotational symmetry breaking of the nanoPET predictions. [81, 82] In Section S2 of the supporting information we make a detailed assessment of the residual symmetry errors of the ShiftML3 model.
When performing hyperparameter optimization we observed that the prediction accuracies of individual nanoPET models do not vary strongly between reasonable guesses of potential hyperparameter choices (cutoffs, number of message passing layers and token size). We therefore employ a set of hyperparameters for the nanoPET implementation that provides a good compromise between accuracy and evaluation speed. An in-depth evaluation of the computational speed of the ShiftML3 model can be found in the section S3 of the supporting information. We choose a radial cutoff of 5 Å to encode local information (displacement vectors, atom types and messages from neighbouring atoms) into tokens. ShiftML3 employs 2 message passing layers and 2 attention layers per message passing step, and a token dimensionality of 192. For better usability, we make ShiftML3 available through a Python package (more information in section S4 of the supporting information).
The predictions are affected by a structure-dependent uncertainty, which should be quantified to assess how reliably the model can be used for a structure determination task – especially to identify the degradation of accuracy that usually occurs in the extrapolative regime. In many cases the uncertainty can be attributed to lack of knowledge of the model due to improper sampling of the training data (epistemic uncertainties). [83] Here, we refer to the prediction accuracy of the model with respect to the DFT calculations, whereas the underlying error of the electronic-structure reference would need to be assessed by comparison with experiments or higher-level calculations.
PET and most other atomistic neural networks make only point estimates of the target quantity. Constructing an ensemble is a universal strategy to equip ML models with uncertainty estimates, [84, 85] even if the model does not allow for a probabilistic interpretation of its predictions. An ensemble model is constructed by training multiple models. The final prediction for an atomic local environment, ${A}_{i}$, of an ensemble of Nens models, is then taken to be the mean of the individual predictions of the “committee” members, $\bar{y}(A_{i})$, and their spread is taken to be an indicator of the reliability of the model predictions:
$\sigma^{2}_{\text{pred}}(A_{i})=\frac{1}{N_{\text{ens}}-1}\sum_{k=1}^{N_{\text{ens}}}\big{[}y^{k}(A_{i})-\overline{y}(A_{i})\big{]}^{2}$ (1)
We compute uncertainty estimates from the committee of PET model predictions. We note in passing that it is often possible to reduce the overhead of ensemble models by building shallow ensembles that share the time-consuming part of the evaluation, and differ only by their last-layer features. [86] Here we opt for a full ensembling to maximise prediction accuracies in line with previous works on prediction chemical shieldings with deep learning networks. [46, 87] Following standard practice, [88, 89] we calibrate the uncertainty estimates of ShiftML3 on the validation set, scaling the predicted ensemble uncertainties with a global, input-independent correction. [57] More details of the model calibration and evaluation of the uncertainty estimates can be found in section S5 of the supporting information.
Figure 1: Parity plots between DFT-computed shieldings and predictions using ShiftML2 or ShiftML3 for (a,b) 1H, (c,d) 13C and (e,f) 15N evaluated on the new CSD-test set. The black line shows perfect correlation while the colour each point is correlated to the prediction uncertainty of that atomic site. Each point is colored according to the rescaled prediction uncertainty associated in each atomic site.
To evaluate the performance of ShiftML3, we first compare predicted isotropic chemical shieldings against DFT calculated isotropic shieldings on the CSD-test set. Figure 1 shows parity plots illustrating the overall agreement between ShiftML3 and DFT with an RMSE of 0.43 ppm, 2.3 ppm and 10.4 ppm for 1H, 13C and 15N, respectively. These correspond to an error reduction between 20% to a factor of 2 relative to ShiftML2. Similar levels of improvement are seen also for almost all nuclei represented in the dataset, as shown in Table 1 (with the exception of 39K, for which there are only 6 atomic enviroments in the test set, making the assessment inconclusive).
Table 1: Prediction accuracies of ShiftML2 (sML2) and ShiftML3 (sML3) models on the CSD-test dataset. MAE and RMSE (in ppm) are evaluated on a hold-out test set.
nucleus MAE RMSE $R^{2}$
sML2 sML3 sML2 sML3 sML2 sML3
${}^{1}\text{H}$ 0.39 0.33 0.51 0.43 0.97 0.98
${}^{13}\text{C}$ 3.15 1.58 4.63 2.32 0.99 1.00
${}^{15}\text{N}$ 10.29 5.00 15.72 10.41 0.98 0.99
${}^{17}\text{O}$ 16.11 7.51 22.96 11.45 0.98 1.00
${}^{19}\text{F}$ 7.77 4.42 10.83 6.66 0.97 0.99
${}^{33}\text{S}$ 34.35 15.05 53.34 26.34 0.88 0.97
${}^{31}\text{P}$ 21.83 9.49 39.20 18.98 0.62 0.91
${}^{35}\text{Cl}$ 17.40 10.44 23.85 15.11 0.96 0.98
${}^{23}\text{Na}$ 7.64 4.09 8.10 4.61 0.00 0.68
${}^{43}\text{Ca}$ 5.81 2.96 6.98 3.34 0.81 0.96
${}^{25}\text{Mg}$ 10.69 3.68 14.55 5.49 0.66 0.95
${}^{39}\text{K}$ 3.42 4.29 4.28 4.65 0.77 0.73
Table 2: Prediction accuracies of ShiftML3 on symmetric tensor components $\sigma_{ij}$ and principal tensor components $\sigma_{\text{PAS}}$. MAE and RMSE (in ppm) are evaluated on a hold-out test set.
nucleus $\sigma_{ij}$ $\sigma_{\text{PAS}}$
MAE RMSE $R^{2}$ MAE RMSE $R^{2}$
${}^{1}\text{H}$ 0.54 0.72 1.00 0.65 0.85 0.98
${}^{13}\text{C}$ 2.10 3.08 1.00 2.64 3.88 1.00
${}^{15}\text{N}$ 5.32 10.75 0.99 7.21 15.62 0.99
${}^{17}\text{O}$ 7.58 12.09 0.99 10.28 16.79 1.00
${}^{19}\text{F}$ 4.05 6.28 1.00 5.73 8.65 0.99
${}^{33}\text{S}$ 17.15 30.37 0.97 22.53 42.73 0.97
${}^{31}\text{P}$ 10.30 18.73 0.98 13.35 23.59 0.95
${}^{35}\text{Cl}$ 10.14 15.32 1.00 14.02 20.50 0.99
${}^{23}\text{Na}$ 2.83 4.01 1.00 4.28 5.24 0.91
${}^{43}\text{Ca}$ 3.14 3.85 1.00 4.53 4.93 0.98
${}^{25}\text{Mg}$ 7.48 9.79 1.00 6.27 8.34 0.93
${}^{39}\text{K}$ 2.93 3.95 1.00 4.42 5.21 0.87
Figure 2: Parity plots between experimental chemical shifts and predictions from ShiftML2, ShiftML3 and DFT calculations for (a,b,c) 1H, (d, e, f) 13C, and (g, h, i)15N evaluated on the benchmark structures from Ramos et al. [14] for 13C and 15N, and the 13 structures where ShiftML2 was evaluated for 1H. The black line shows the linear regression between experimental chemical shifts and predicted isotropic shiedings with their slope and intercept indicated in each respective subplot.
We then evaluate the chemical shielding tensor predictions of the ShiftML3 model on the CSD-test set, and report the prediction accuracies of symmetric tensor components in Table 2. More information, including parity plots for 13C and 15N anisotropic shieldings can be found in section S5 of the supporting information. We report prediction accuracies for the combined principal components $\sigma_{\text{PAS}}$ and combined tensor components $\sigma_{ij}$. We report averaged prediction errors over all tensor components or principal components and across atomic environments in the test dataset, consistent with the error metric reported for chemical shieldings, which themselves are the average of the three principal components of the chemical shielding tensor. We obtain averaged eigenvalues of the ensemble, by first computing the eigendecomposition of the individual committee member predictions and then averaging over the committee prediction of principal components.
Figure 3: Parity plots between experimental principal components $\sigma_{PAS}$ of the CSA tensor and predicted components using ShiftML3 for the CSA components evaluated on a subset of benchmark structures from Ramos et al. [14] The black line shows the linear regression between experimental and predicted CSA principal components, with their slope and intercept shown in each respective subplot.
The improvements in prediction accuracy against the DFT reference also translate to an increased accuracy of ShiftML predictions against experiments. We find that on a representative benchmark set of experimentally measured 1H and 13C shieldings, [49, 14] ShiftML3 predicts chemical shieldings with accuracies almost on par with GIPAW DFT, namely: 0.53(0.49) ppm for 1H and 2.4(2.3) ppm for 13C, for ShiftML3(DFT). Figure 2, shows parity plots of isotropic chemical shieldings estimated with ShiftML2, ShiftML3 and DFT with respect to experimental chemical shifts. DFT values were evaluated recomputing the NMR shieldings using GIPAW [7] on the the Quantum ESPRESSO software [90, 91] with the same DFT parameters as in Ref. 49. The conversion between experimental chemical shifts and predicted shieldings is performed by fitting globally the shieldings across all test compounds to a linear form, also shown in figure 2. Further details about the experimental benchmark are discussed in section S7 of the supporting information.
We also assess the accuracy of the CSA predictions against experiments, based on a benchmark of 13C shieldings for a set of structures whose experimental shifts had been reported in Ref. 72. The accuracy in terms of the three principal components ($\sigma_{11},\sigma_{22},\sigma_{33}$) of the shielding tensor is on par with that of DFT. The RMSE of for ShiftML3 (5.2 ppm) is even smaller than for DFT (5.8 ppm) which is likely fortuitous (Fig. 3).
Overall, we find that the improved accuracy of ShiftML3 against DFT-computed isotropic shieldings, relative to ShiftML2, is reflected in a corresponding improvement of the match with experiments, that is now essentially limited by the accuracy of the reference DFT calculations. The tensorial predictions also have a DFT-level match with experiments, providing a further diagnostic to support structure determination and provide mechanistic insights.
To summarize our findings, we have presented a machine learning model for the prediction of chemical shieldings in organic crystals, using an unconstrained graph neural network architecture that delivers chemical shielding tensors and uncertainty quantification through an ensemble structure. We have shown that across all nuclei that are well-represented in the dataset, ShiftML3 predicts accurate shielding tensors and isotropic shieldings in organic crystals. The isotropic prediction RMSE of 0.43 ppm for 1H and 2.3 ppm for 13C against GIPAW reference calculations are a major improvement over previous ShiftML versions based on symmetry-adapted kernel models. The improved prediction accuracies translate to an accuracy against experimental benchmarks that practically matches that of DFT calculations, improving the confidence of NMR-guided structural determination. The availability of chemical shieldings anisotropy, with an accuracy that also approaches that of DFT, facilitates comparison with other types of NMR experiments, providing structural and mechanistic insights. The easy-to-use, fast, and openly-available implementation will facilitate usage for a broad community interested in the theoretical and experimental study of organic materials by means of NMR.
Acknowledgements.
.1 ACKNOWLEDGEMENTS
We thank Filippo Bigi and Michelangelo Domina for assistance with their nanoPET-IST implementation and Sergey Pozdnyakov for adapting the original PET architecture for isotropic shieldings. We thank Joseph William Abbott for the provision of a $\lambda$-SOAP model, which we used to protoype chemical shielding anisotropy predictions. This work has been supported by Swiss National Science Foundation Grant No. 200020_212046, and by the NCCR MARVEL, funded by the Swiss National Science Foundation (grant number 182892). MC also acknowledged funding from the ERC Horizon 2020 Grant No. 101001890-FIAMMA.
.2 SUPPORTING INFORMATION
The supporting information is provided below, containing additional details on dataset and model construction, uncertainty quantification, and the usage of the ShiftML3 Python package, which is available under an open source license ad https://github.com/lab-cosmo/shiftml. The training, validation and testing datasets containing chemical shielding tensor data as well as training input files, model files and evaluation scripts will be made available upon publication via the public data record associated with this publication. [71]
Supplementary Information
S1 Detailed prediction accuracies of ShiftML3
We evaluate the accuracy of ShiftML3 on the CSD-test set described in the main text. In Table S1, we report the size of the training, validation and CSD-test set together with the mean average error (MAE), root-mean-squared-error (RMSE), the coefficent of determination (R2 ), and the standardized RMSE ( % ) evaluation metrics of $\sigma_{\text{iso}}$ predictions of ShiftML3 for all chemical species contained in CSD-test.
Table S1: Predicted accuracy of $\sigma_{\text{iso}}$ of ShiftML2 (sML2) and ShiftML3 (sML3) on the CSD-test hold-out set
[…]
sML2 sML3 sML2 sML3 sML2 sML3 sML2 sML3
1H 547,887 18,505 47,706 0.39 0.33 0.51 0.43 16 14 0.97 0.98
13C 437,529 16,971 43,435 3.15 1.58 4.63 2.32 9 5 0.99 1.00
15N 119,413 1,879 4,635 10.29 5.00 15.72 10.41 14 10 0.98 0.99
17O 135,305 3,272 8,058 16.11 7.51 22.96 11.45 13 6 0.98 1.00
19F 37,765 296 569 7.77 4.42 10.83 6.66 18 11 0.97 0.99
33S 30,500 386 1,084 34.35 15.05 53.34 26.34 35 17 0.88 0.97
31P 8,508 84 151 21.83 9.49 39.20 18.98 61 30 0.62 0.91
35Cl 25,608 188 569 17.40 10.44 23.85 15.11 21 13 0.96 0.98
23Na 1,138 9 5 7.64 4.09 8.10 4.61 100 57 0.00 0.68
43Ca 613 3 5 5.81 2.96 6.98 3.34 44 21 0.81 0.96
25Mg 283 3 7 10.69 3.68 14.55 5.49 58 22 0.66 0.95
39K 995 3 6 3.42 4.29 4.28 4.65 48 52 0.77 0.73
S2 Rotational fluctuations of PET and PET-ensemble
PET models learn the correct rotational behaviour of the targets via rotational augmentations. Langer and coworkers have investigated the effect relaxing exact rotational invariance to learned invariances, [74] and found that for realistic atomistic modelling tasks, this constraint can be safely relaxed. The remaining rotational fluctuations of the PET model and PET ensemble should however be carefully assessed for chemical shielding prediction tasks. Unlike the prediction of potential energies, or other structure wise global quantities, chemical shieldings are atom-wise properties. The common ansatz of predicting potential energies with neural networks, is by summing over local, per-atom terms. Shortcomings (imperfectly learned rotational invariance) of a non-invariant architecture might simply be compensated by error cancellation summing over atomic contributions. Since chemical shieldings and shielding tensors are exactly these local contribution (one prediction per atom), we study the effects of applying random rotations to structures and measuring the fluctuations of the predicted isotropic chemical shieldings. We apply 30 random rotations to each structure in the CSD test set and evaluate the standard deviation of the isotropic chemical shielding predictions. For a perfectly invariant model, these fluctuations should be exactly zero due to the invariance of the model. For an approximately invariant model, these predictions should be quite small, at most a fraction of the prediction RMSEs, such that they do not interfere with NMR crystal structure prediction workflows. In Table S2 we list rotational fluctuations, of ShiftML3 predictions across various chemical species and compare them with the rotational fluctuations of single PET models, highlighting the strongly reduced fluctuation of the ensemble of PET models that make ShiftML3.
[…]
nucleus single ensemble (ShiftML3)
[…]
${}^{1}\text{H}$ 0.41 1.78 0.06 0.23
${}^{13}\text{C}$ 1.19 7.13 0.14 0.71
${}^{15}\text{N}$ 2.57 23.49 0.29 1.57
${}^{17}\text{O}$ 3.98 43.35 0.45 2.22
${}^{19}\text{F}$ 2.27 9.09 0.26 0.70
${}^{33}\text{S}$ 5.28 36.90 0.59 3.13
${}^{31}\text{P}$ 2.82 19.81 0.33 1.49
${}^{35}\text{Cl}$ 4.73 22.13 0.50 2.03
We find that for most nuclei, ensemble uncertainty estimates are correlated with the rotational fluctuations.
[…]
We evaluate the structure wise fluctuations of the RMSEs of the ShiftML3 predictions for CSD-test with respect to the GIPAW reference. To assess the effect of approximate invariance we also compute the fluctuations of the RMSE between all 1H environments in a structure and it non rotated structure. The mean fluctuation (mean of the standard deviation of fluctuations) of this RMSE across the CSD test set is 0.0065 ppm and the maximal fluctuation we find to be 0.0271 ppm.
S3 Model evaluation timings
We benchmark the evaluation time of the ShiftML3 model. All benchmark timings are obtained via the ASE interface of the ShiftML library. We report timings that include setting up the ASE calculator once and then computing chemical shielding tensors for a series of structures. We believe that this protocol reflects most accurately the evaluation timings for a common NMR crystallography workflow. The timings via the ASE interface include also the neighbour list construction. The ASE interface does not support batched evaluations, or efficient parallelization over ensemble models for smaller system sizes. We report CPU-only and CPU+GPU timings. We report evaluation times on a Lenovo workstation with a AMD Ryzen Threadripper PRO 3955WX CPU with 16 cores. CPU+GPU timings are obtained on the same workstation, using a NVIDIA GeForce RTX 4070 Ti-super consumer GPU, which is representative of a configuration that does not require access to high-end AI-dedicated hardware. As shown in Table S3, even a low-end GPU accelerates by 4x-6x the evaluation of structures with the size of typical small-molecule pharmaceuticals. The acceleration is also evident at larger system sizes: the scaling of the ShiftML3 model for diamond supercells of increasing size is almost perfectly linear, and there is an order-of-magnitude acceleration when using a GPU (Fig. S2). These timings are also representative of large biomolecules, as hinted at by the inference timings for the protein crystal of 1PGA [92], the B1 immunoflobulin-binding domain of the streptococcal protein G with 56 amino acid residues and 3,656 atoms in the conventional cell (also shown in Fig. S2). 1PGA was a target of several solid-state NMR studies [93, 94] and therefore lies well within the range of atomistic systems of interest for which chemical shielding predictions may complement experimental studies. A computation of chemical shieldings for this protein takes 5.6 s using the GPU. Moving to even larger systems, we find that on the consumer GPU used for benchmarking with 16GB of VRAM, we can predict 13C shieldings in diamond supercells of up to 16,000 atoms before exhausting the graphics card video memory, demonstrating the memory efficiency of ShiftML3.
Table S3: Average timing of the ShiftML3 model for evaluating chemical shieldings on CPU or GPU. We also list the average number of atoms $\overline{\text{N}}_{\text{atoms}}$ in the CSD-test hold-out test set.
Set Time per struct. [s] CPU/GPU $\overline{\text{N}}_{\text{atoms}}$
CSD-test 0.81/0.23 125
Figure S2: Scaling of the inference time of the ShiftML3 model with system size. Blue lines indicate timings for diamond supercells of increasing size. Black markers are the evaluation times of the 1PGA protein crystal. Grey lines correspond to linear scaling.
S4 Implementation
To simplify usage of ShiftML3 for inference, we make it available with an easy-to-install Python package, released together with this publication. ShiftML is available from the Python Package Index PyPi and can be easily installed with
 pip install shiftml
The ShiftML3 model integrates into existing materials modelling and crystal structure prediction workflows via the atomic simulation environment. Chemical shielding predictions can be obtained by initializing an ASE calculator object and reading in the geometry file in one of the various supported file formats. Once the calculator is initialized, the model weights are automatically downloaded and loaded by the model. A minimalistic usage example is given below, and a more thorough tutorial is available at http://atomistic-cookbook.org/shiftml/shiftml-example.html.
from ase.build import bulk
from shiftml.ase import ShiftML
frame = bulk("C", "diamond", a=3.566)
model = ShiftML("ShiftML3")
Ypred = model.get_cs_tensor(frame)
S5 Uncertainty quantification and calibration
The ensemble uncertainty estimates we obtain are miscalibrated (as it is often the case [88, 89] in uncertainty quantification schemes for neural networks), i.e. the uncertainty estimates across all samples of a hold-out validation or test set model uncertainty estimates are systematically over- or underconfident. It is possible to rescale the uncertainty estimates, and correct the miscalibration by computing a global, input-independent scaling factor [57, 95] (an empirical correction that is similar in spirit to temperature scaling[88] a classical method in uncertainty quantification). The scaling factor can be computed via a closed form expression, that can be derived by maximizing the Gaussian negative log likelihood on a hold out test set given predicted uncertainties and squared prediction errors $z^{2}_{i}(A)$.
$\alpha^{2}=\sum^{N_{\text{samples}}}_{i=1}\frac{z^{2}_{i}(A)}{\sigma^{2}_{i,\text{pred}}(A)}$ (2)
The final, calibrated, uncertainties are then obtained as $\alpha^{2}\sigma^{2}_{\text{pred}}(A)$ although in practice we scale the ensemble members around their mean, which simplifies uncertainty propagation. [86] We find that for isotropic chemical shieldings and shielding tensor principal components the calibration factors $\alpha$ are moderate and always close to $1$, indicating that the ShiftML3 ensemble uncertainty estimates, are already well calibrated.
We evaluate the uncertainty estimates of the ShiftML3 model on the CSD-test set and plot the predicted uncertainties of 13C shieldings vs the actual prediction errors. In Fig. S3 we show parity plots of predicted and calibrated uncertainties of the ShiftML3 model and the actual errors on a hold out test set.
Figure S3: Predicted uncertainties of 13C isotropic chemical shieldings $\text{std}_{\text{pred}}[\sigma_{\text{iso}}]$ vs the actual prediction errors $\text{z}[\sigma_{\text{iso}}]$ (a) and predicted uncertainties of the corresponding principal components $\text{std}_{\text{pred}}[\sigma_{\text{PAS}}]$ vs the actual component prediction errors $\text{z}[\sigma_{\text{PAS}}]$ (b). The grey lines indicate quantiles of the expected error distribution for a given predicted uncertainty values and are intended to guide the readers eye: For example, on the test set, for well calibrated uncertainty estimates there should be 90% of the samples between the two outmost lines as they correspond to quantile lines of 95% and 5% of the folded log normal distribution.
S6 Antisymmetric chemical shielding tensor predictions
The chemical shielding tensor can be decomposed into a symmetric and antisymmetric part (eq. 3):
$\sigma=\sigma_{\text{sym}}+\sigma_{\text{antisym}}=\frac{1}{2}(\sigma+\sigma^{\text{T}})+\frac{1}{2}(\sigma-\sigma^{\text{T}})$ (3)
Only the symmetric part enters directly into the NMR spectrum, whilst the antisymmetric part effects relaxation times. [96] Therefore, in the main part of this manuscript we have focussed on evaluating and utilizing the CSA, $\sigma_{\text{PAS}}$ which are the principal components of $\sigma_{\text{sym}}$. ShiftML3 can predict the full chemical shielding tensor, containing symmetric and antisymmetric parts. In Figure S4 we show parity plots of the symmetric and antisymmetric components $\sigma_{ij,\text{sym}}$ and $\sigma_{ij,\text{antisym}}$ of the 13C and 17O chemical shielding tensors.
Figure S4: Parity plots of symmetric and antisymmetric chemical shielding tensor components $\sigma_{ij,\text{sym}}$, $\sigma_{ij,\text{antisym}}$ for 13C (a,b) and 17O (c,d) comparing the values predicted with ShiftML3 and those computed with DFT.
In Table S4 we list the prediction accuracies of ShiftML on CSD-test separately on antisymmetric and symmetric components of the chemical shielding tensor.
Table S4: Prediction accuracies RMSEs (ppm) and normalized prediction accuracies RMSE/std (%) of ShiftML3 of the components of the symmetric $\sigma_{ij,\text{sym}}$ and antisymmetric shielding tensor $\sigma_{ij,\text{antisym}}$.
nucleus $\sigma_{ij,\text{sym}}$ $\sigma_{ij,\text{antisym}}$
[…]
${}^{1}\text{H}$ 0.72 6 0.52 57
${}^{13}\text{C}$ 3.08 5 2.28 41
${}^{15}\text{N}$ 10.75 11 5.98 47
${}^{17}\text{O}$ 12.09 7 7.82 27
${}^{19}\text{F}$ 6.28 5 3.35 39
${}^{33}\text{S}$ 30.37 18 18.10 52
${}^{31}\text{P}$ 18.73 14 11.02 72
${}^{35}\text{Cl}$ 15.32 5 9.57 47
${}^{23}\text{Na}$ 4.01 2 1.58 81
${}^{43}\text{Ca}$ 3.85 1 1.31 29
${}^{25}\text{Mg}$ 9.79 3 4.55 39
${}^{39}\text{K}$ 3.95 1 2.05 67
Figure S5 shows that the symmetric and antisymmetric tensor components (for nuclei like 13C and 17O) have similar absolute prediction errors. However, because the antisymmetric values span a much smaller range, their relative errors are much larger than those of the symmetric components.
Figure S5: Parity plots of symmetric and antisymmetric chemical shielding tensor components $\sigma_{ij,\text{sym}}$, $\sigma_{ij,\text{antisym}}$) for 13C (a) and 17O (b) - combined parity plots.
In Figure. S6 we show parity plots of symmetric shielding tensor component predictions and principal tensor component predictions.
Figure S6: Parity plots of principal tensors components $\sigma_{\text{PAS}}$ and symmetric tensor components $\sigma_{ij}$ for 13C (a,b) and 15N (c,d)
S7 Details of experimental benchmark evaluations
Evaluation of 1H was done on the same 13 structures where ShiftML2 was evaluated [49] on the same geometry optimized structures. Predicted shieldings of methyl or primary amine groups were averaged out accounting for rapid rotation, resulting in the averaged sites have equivalent chemical shielding. Ambiguous experimental assignments were solved by iterating along all possible permutations of those unassigned shifts and taking the assignment with the lowest RMSE between predicted and experimental chemical shift such as for methylene groups. Only the permutation with the lowest RMSE entered the global linear regression. We note that the reported RMSE comparing prediction to experiment for ShiftML2 is modestly different than the original work. This difference originates from the use of a global regression as compared to on-the-fly regression used in the original work. For 13C and 15N, the accuracy was were evaluated from the benchmark consisting of of 36 structures from Ramos et al. [14] on geometries optimized on the PBE-D3(BJ)level of theory [97]. For the carbon CSA parameters, we used the benchmark consisting of 19 structures from Hartman et al. [72]
Table S5: RMSE and Linear regression parameters are for the DFT, ShiftML2, and ShiftML3 prediction and experimental chemical shifts from the respective benchmark.
Nucleus Method RMSE Slope Intercept
1H iso ShiftML2 0.59 -0.8985 27.86
ShiftML3 0.53 -0.9024 28.05
DFT 0.49 -0.8902 27.58
13C iso ShiftML2 2.91 -0.9667 165.07
ShiftML3 2.44 -0.9732 166.23
DFT 2.34 -0.9736 166.06
15N iso ShiftML2 14.03 -1.0556 184.82
ShiftML3 7.24 -1.025 183.34
DFT 5.85 -1.0137 181.77
13C PAS ShiftML3 5.2 -0.9576 164.84
DFT 5.85 -0.9525 164.27
S8 Comparing ShiftML3 with other machine learning models for chemical shielding prediction
A direct comparsion of the ShiftML3 prediction accuracies with ShiftML2 reveals that it is more accurate for predicting isotropic chemical shieldings across all relevant elements. To the best of our knowledge, no other machine learning model has been trained on the last iteration of the ShiftML datasets, including thermally distorted structures and additional chemical elements. Unzueta et al. trained GNN models [87] and Liu and coworkers a multiresolution 3D-DenseNet architecture on the CSD-2K dataset [46], an early iteration of the ShiftML training sets. An alternative model construction strategy is the NMRNet from Xu and coworkers, that finetuned a transformer model for the chemical shielding prediction in liquids on the CSD-2K set [98]. All three works report evaluation accuracies on the CSD-500 set, a set of DFT relaxed organic crystals, containing at most H,C,N and O. Unfortunately the number of shared structures between the CSD-test and the CSD-500 test set is too small (about 10 structures) to make direct comparisons between the accuracies of ShiftML3 and other neural network architectures. Shielding computations are highly sensitive to k-space discretization and choice of pseudopotentials [99]. CSD-2K/CSD-500 was computed with other DFT convergence parameters than the ShiftML2/ShiftML3 dataset, hampering severely a direct comparison between all models - we regularly find that shielding values from computations employing different, yet reasonable convergence parameters, differ (in terms of their RMSEs) nearly as much as the ShiftML3 predicted shieldings differ from the CSD test reference calculations. Instead, we evaluate ShiftML2 and ShiftML3 on 351 structures in our test set that only contain H,C,N,O and are not MD distorted structures (containing 7536 1H, 6546 13C, 665 15N and 1311 17O test environments) and compare their relative improvements in accuracies against the other models, to give some context to the gains in accuracy we achieve with ShiftML3. In Table S6, we compare prediction accuracies of the original ShiftML1.0 model, Unzuetas ensemble of GNNs, Liu’s MR-3D-DenseNet and compare them with the prediction accuracies of ShiftML2 and ShiftML3 on the H,C,N,O-relaxed subselection of the CSD-test set.
Table S6: Test set RMSEs (ppm) of the original ShiftML1.0 model (KRR [45]), Unzuetas GNNs (GNN [87]) and Liu’s MR-3D-DenseNet (DenseNet [46]) and Xu’s NMRNet (NMRNet [98]) on the CSD-500 test set. We compare these values with the prediction accuracies of ShiftML2 (sML2) and ShiftML3 (sML3) on a comparable subset of the CSD test set. Note that KRR, GNN, Densenet and NMRNet are evaluated on the CSD-500 test set. As discussed in the text, due to the small overlap of CSD-500 and the CSD test set, we choose to evaluate sML2 and sML3 on a subset of the CSD test set containing only H,C,N,O structures that are geometry relaxed. (*) Denotes, that the results of sML2 and sML3 were obtained on that subselection.
Nucleus KRR GNN DenseNet NMRNet sML2* sML3*
1H 0.49 0.49 0.37 0.35 0.47 0.39
13C 4.3 4.06 3.3 3.21 4.07 1.97
15N 13.3 9.90 10.2 9.45 12.52 5.71
17O 17.7 14.4 15.3 13.03 19.50 9.91